# Discovery of high-frequency quasi-periodic oscillation in short-duration gamma-ray bursts

### v1

@@ Abstract
Rapidly rotating newborn magnetars, which originate from binary neutron star (NS) mergers and serve as the central engines of short gamma-ray bursts (GRBs), may leave some imprints on their prompt gamma-ray lightcurves even though they are far from their radiating fireballs. A high-frequency quasi-periodic oscillation (QPO) would be a unique feature for the magnetar central engine, especially a hypermassive magnetar. By conducting a systematic analysis of the prompt gamma-ray lightcurves from 605 short GRBs observed by Fermi/GBM, we have identified such QPO signals in three GRBs (e.g., GRB 120323A, GRB 181222B, and GRB 190606A). The QPOs that peaked at $1258^{+6}_{-6}$ Hz for GRB 120323A, $623^{+4}_{-4}$ Hz for GRB 181222B, and $1410^{+4}_{-5}$ Hz for GRB 190606A are all with a confidence level above 5.2 $\sigma$. The high-frequency QPO signals of those three short GRBs may be caused by a hypermassive magnetar acting as the central engine in a binary neutron star merger of binary neutron star.
@@ 1 Introduction
Since the discovery of gamma-ray bursts (GRBs) in 1963, several lines of observational evidence suggest that the progenitors of short-duration GRBs (Kumar & Zhang, 2015) or nearby short-duration GRBs with extended emission (Rastinejad et al., 2022; Yang et al., 2022; Troja et al., 2022; Chang et al., 2023; Gompertz et al., 2023; Lü et al., 2022; Ferro et al., 2023; Du et al., 2024a, b) originate in the mergers of two compact stellar objects. The leading model is neutron star-neutron star mergers (NS-NS)(Paczynski & Anderson, 1986; Eichler et al., 1989) or neutron star-black hole mergers (NS-BH) (Paczynski, 1991). The first direct detection of the NS-NS merger event GW 170817 that was associated with short GRB 170817A and kilonova AT2017gfo, has firmly verified the NS-NS merger model at least for some short GRBs (Abbott et al., 2017; Goldstein et al., 2017; Savchenko et al., 2017; Zhang et al., 2018; Lü et al., 2019).
@@
Many groups have studied the mergers of two NSs in numerical relativity simulations (Rosswog et al., 2003; Rezzolla et al., 2011; Hotokezaka et al., 2013; Nagakura et al., 2014), but the origin of the remnants of NS-NS mergers remains an open question (Zhang, 2011). Since both the NS equation of state and the nascent NS mass are poorly constrained, there are four types of merger remnants that may be formed under certain conditions (Rosswog et al., 2000; Dai et al., 2006; Fan & Xu, 2006; Rezzolla et al., 2010; Giacomazzo & Perna, 2013; Zhang, 2013; Lasky et al., 2014; Rosswog et al., 2014; Gao et al., 2016; Li et al., 2016; Lü et al., 2018; Sarin & Lasky, 2021): (1) a black hole surrounded by a dense torus; (2) a differential-rotation-supported hypermassive NS (HMNS) which may survive for $\sim$300 ms before collapsing into a BH (Rosswog et al., 2003; Paschalidis et al., 2012); (3) a supramassive NS (SMNS) supported by rigid rotation, which may survive for hundreds of seconds before collapsing to a BH (Rowlinson et al., 2010; Lü et al., 2017, 2020); (4) a stable NS with a much longer lifetime (Dai et al., 2006; Zhang, 2013; Lü et al., 2015; Gao et al., 2016). The dynamical nature of the NS-NS merger suggests that quasi-periodic signals, quasi-periodic oscillations (QPOs), are more probable if the central engine is a HMNS (Chirenti et al., 2019), or if internal plateau [^1] A fairly constant emission followed by a steep decay (Troja et al., 2007). emission (or plateau emission) in X-rays could be shown at a later time within the SMNS (or stable NS) as the central engine (Dai & Lu, 1998; Zhang & Mészáros, 2001).
@@
From an observational point of view, a small fraction of short GRBs that have plateau (or internal plateau) emission in X-rays point toward a stable NS (or a SMNS) as the central engine of short GRBs (Dai & Lu, 1998; Zhang & Mészáros, 2001; Rowlinson et al., 2013, 2010; Lü et al., 2015, 2017). However, due to their short lifetimes it is difficult to observe definite clues for the HMNS as the central engine of short GRBs. Recently, Chirenti et al. (2023) discovered kilohertz quasi-periodic oscillations in two short GRBs with a confidence level about $5\sigma$ from archival Burst and Transient Source Experiment (BATSE) data, and suggest that these are evidence for a shortly surviving HMNS before the final collapse into a black hole.
@@
Motivated by the above described search in the BATSE data for a QPO that is the signature of a HMNS, this paper presents a systematic analysis of the short-duration GRBs observed with Fermi/GBM since its inception in 2008, and investigates whether such QPOs are also hidden in the light curves of those short GRBs. We systematically searched more than 3674 GRBs observed by Fermi/GBM, and found 605 short-duration GRBs. We focused on analysing the 605 short-duration GRBs, and found that three short-duration GRBs (e.g., 120323A, 181222B, and 190606A) with a high-frequency quasi-periodic signal can be identified with a confidence level above 5.2 $\sigma$. Our data reduction and QPO search method are shown in §2. The observations and results of our QPO search for those three GRBs are presented in §3. Conclusions are drawn in §4 with some additional discussions.
@@ 2.1 Fermi/GBM data reduction
The Fermi satellite was launched in June 2008 and has been in operation for 16 years. There are two instruments onboard the Fermi satellite. One is the Gamma-ray Burst Monitor (GBM)(Meegan et al., 2009), which has 12 sodium iodide (NaI) and two bismuth germanate (BGO) scintillation detectors covering the 8 keV-40 MeV energy band. The other is the Large Area Telescope (LAT)(Atwood et al., 2009), which has an energy coverage from 20 MeV to 300 GeV.
@@
We find that the $T_{\rm 90}$ of 605 GRBs out of 3674 GRBs are less than 2 seconds at the official Fermi web site. Here, we focus on studying the 605 short-duration GRBs, and extract the light curves of those short-duration GRBs from 12 NaI detectors by adopting a 128 $\mu$s time bin.
@@ 2.2 QPO search method in short GRB samples
The data that we adopted consist of flux time series of 605 short GRBs, and it includes 12 simultaneous light curves from the 12 NaI detectors in the GBM TTE data for each short GRB. We extracted the NaI detector light curve with a time-bin of 128 $\mu$s in the energy range of 50-900 keV. Each NaI detector points to a slightly different position on the sky, the light curves of each GRB are ranked in order of increasing angular distance between the GRB position and detector pointing. They are designated as $LC_{i}$ where $i=1,2...,12$ from closest to farthest. Twelve summed light curves are constructed by successively adding one more light curve to the sum in order of the ranking. These are designated as
@@
These summed light curves are not statistically independent. Next, we applied to $LC_{\rm sum}(n)$ and cut down the light curves to a duration of 2.048 seconds, which consists of 16000 bins. Our selection of the start and end times of the light curves can be expressed as
@@
Here, the $a_{\rm j}$ is the result of the Discrete Fourier transform, and $N$ is the number of bins in a light curve. Each power spectrum $I_{\rm j}$ is fitted with a broken power-law plus constant using the maximum likelihood method (Vaughan, 2010) by assuming that the powers are exponentially ($\chi^{2}_{v}$ with 2 dof) distributed. The broken power-law function can be expressed as
@@
where $\alpha_{1}$ and $\alpha_{2}$ are the exponential slopes, respectively. $j_{\rm b}$ is the break point between the two slopes, and $C$ is the white noise component of the power spectrum. The likelihood function is equivalent to minimizing the function $D$ which is expressed as
@@
$R_{\rm k-max}$ is the largest value in the power spectrum $R_{\rm k}(j)$. Each GRB contains 12 light curves, and we search each power spectrum for 10 values of $k$ and therefore obtain 120 $R_{\rm k-max}$ values for each GRB. In order to evaluate whether or not $R_{\rm k-max}$ is generated by quasi-periodic signals, one can calculate the probability that $R_{\rm k-max}$ is not produced by noise. A simulation is performed generating $10^{10}$ power spectra from $10^{10}$ simulated Poisson noise light curves. For each simulated power spectrum, ten $R_{\rm k-max}$ ($k$=1,…,10) values are generated by the method described above. To estimate a confidence level associated with a given $R_{\rm k-max}$, the number of times $N_{\rm k-max}$ that this $R_{\rm k-max}$ is exceeded in the $10^{10}$ simulated power spectra is counted. The probability $G_{\rm k}$ of a given $R_{\rm k-max}$ value is then calculated as
@@
$G_{\rm k}$ is the probability that $R_{\rm k-max}$ is not generated by random noise. Therefore, for each GRB, we can obtain $120$ different $G_{\rm k}$ values. If there is a quasi-periodic signal in a GRB, the maximum $G_{\rm k}$ value is the value that is most likely to be produced by the quasi-periodic signal. This procedure can be represented as
@@
Here, $S_{j}$ is the best fit with broken power-law function, and it is not a free parameter in the fits. $A$ is the power density of the Lorentzian at its center, $v$ is the central frequency of the Lorentzian, and $\Delta v$ is the full width at half maximum (FWHM) of the Lorentzian. We adopt the central frequency $v$ as the representative frequency of the quasi-periodic signal. By considering the frequency of the quasi-periodic signal which is distributed within a range, we use 90% integral area of the $L_{j}$ function within the Fourier frequency as the representative frequency range of the quasi-periodic signal. The lower limit and upper limit of the frequency correspond to the integral area 5% and 95% , respectively.
@@ 4 Conclusions
Based on the data analysis of our QPO search for those three short GRBs, we find that the duration of a detected QPO is shorter than that of the burst. For example, the start and end time of the burst for GRBs 120323A, 181222B, and 190606A are [0, 0.384] s, [0.032, 0.576] s, and [0, 0.224] s, respectively. However, the start and end time of the detected QPO for those three short GRBs are [0.0384$\pm$0.0256, 0.0896$\pm$0.0256] s, [0.064$\pm$0.0256, 0.1664$\pm$0.0256] s, and [0.01$\pm$0.0256, 0.112$\pm$0.0256] s, respectively. On the other hand, the light curves of those three short GRBs seem to be composed of two pulses, and it is found that the QPO signals exist during the first pulse and the rising segment of second pulse, but disappear during the decay segment of the second pulse. In general, two effects may result in such a situation. One is incorrect data analysis with the selected time step, but we adopted a time step of duration 0.0512 s which is much less than the burst duration. This smaller time step is not likely to result in such a situation. The QPO signal disappearing during the decay segment of the second pulse may arise in the intrinsic physics, but this remains debatable.
@@
However, by invoking a HMNS to interpret the high-frequency QPO in the prompt emission of those three short GRBs (e.g., GRBs 120323A, 181222B, and 190606A), imposes two curious questions that are not clarified in either theory or numerical relativity simulations of neutron star mergers. One question regards the composition and structure of a HMNS. Initially, the temperature should be high enough shortly after the formation of the HMNS, and the composition of the HMNS should be dominated by fluid instead of rigid structure because there has not been enough time to cool down. If this is the case, the oscillation of fluid seems to have difficulty producing such high-frequency quasi-periodic signals. The other question is how to produce the QPO signals only in the first pulse and the rising segment of the second pulse. There remains poor understanding of this question.
In order to investigate the remnants of those three short GRBs-like events and find evidence for a HMNS in a double NS merger, more information from the simultaneous detection of both gravitational wave and multi-band electromagnetic radiation is essential. We therefore strongly encourage follow-up observations when the LIGO-Virgo-KAGRA (LVK) collaboration detects the gravitational wave signal of those three short GRBs-like events in the future.

### v2

@@ Abstract
Rapidly rotating newborn magnetars, which originate from binary neutron star (NS) mergers and serve as the central engines of short gamma-ray bursts (GRBs), may leave some imprints on their prompt gamma-ray light curves even though they are far from their radiating fireballs. A high-frequency quasi-periodic oscillation (QPO) would be a unique feature for the magnetar central engine, especially a hypermassive magnetar. By conducting a systematic analysis of the prompt gamma-ray light curves from 605 short GRBs observed by Fermi/Gamma-ray Burst Monitor, we have identified such QPO signals in three GRBs (e.g. GRB 120323A, GRB 181222B, and GRB 190606A). The QPOs that peaked at $1258^{+6}_{-6}$ Hz for GRB 120323A, $623^{+4}_{-4}$ Hz for GRB 181222B, and $1410^{+4}_{-5}$ Hz for GRB 190606A are all with a confidence level above 5.2 $\sigma$. The high-frequency QPO signals of those three short GRBs may be caused by a hypermassive magnetar acting as the central engine in a binary NS merger of a binary NS.
@@ 1 Introduction
Since the discovery of gamma-ray bursts (GRBs) in 1963, several lines of observational evidence suggest that the progenitors of short-duration GRBs (Kumar & Zhang, 2015) or nearby short-duration GRBs with extended emission (Rastinejad et al., 2022; Yang et al., 2022; Troja et al., 2022; Chang et al., 2023; Gompertz et al., 2023; Lü et al., 2022; Ferro et al., 2023; Du et al., 2024a, b) originate in the mergers of two compact stellar objects. The leading model is neutron star–neutron star (NS–NS) mergers (Paczynski & Anderson, 1986; Eichler et al., 1989) or neutron star–black hole (NS–BH) mergers (Paczynski, 1991). The first direct detection of the NS–NS merger event GW170817, which was associated with short GRB 170817A and kilonova AT2017gfo, has firmly verified the NS–NS merger model at least for some short GRBs (Abbott et al., 2017; Goldstein et al., 2017; Savchenko et al., 2017; Zhang et al., 2018; Lü et al., 2019).
@@
Many groups have studied the mergers of two neutron stars (NSs) in numerical relativity simulations (Rosswog et al., 2003; Rezzolla et al., 2011; Hotokezaka et al., 2013; Nagakura et al., 2014), but the origin of the remnants of NS–NS mergers remains an open question (Zhang, 2011). Since both the NS equation of state and the nascent NS mass are poorly constrained, there are four types of merger remnants that may be formed under certain conditions (Rosswog et al., 2000; Dai et al., 2006; Fan & Xu, 2006; Rezzolla et al., 2010; Giacomazzo & Perna, 2013; Zhang, 2013; Lasky et al., 2014; Rosswog et al., 2014; Gao et al., 2016; Li et al., 2016; Lü et al., 2018; Sarin & Lasky, 2021): (1) a black hole (BH) surrounded by a dense torus; (2) a differential-rotation-supported hypermassive NS (HMNS), which may survive for $\sim$300 ms before collapsing into a BH (Rosswog et al., 2003; Paschalidis et al., 2012); (3) a supramassive NS (SMNS) supported by rigid rotation, which may survive for hundreds of seconds before collapsing to a BH (Rowlinson et al., 2010; Lü et al., 2017, 2020); and (4) a stable NS with a much longer lifetime (Dai et al., 2006; Zhang, 2013; Lü et al., 2015; Gao et al., 2016). The dynamical nature of the NS–NS merger suggests that quasi-periodic signals and quasi-periodic oscillations (QPOs) are more probable if the central engine is an HMNS (Chirenti et al., 2019), or if internal plateau  [^1] A fairly constant emission followed by a steep decay (Troja et al., 2007). emission (or plateau emission) in X-rays could be shown at a later time within the SMNS (or stable NS) as the central engine (Dai & Lu, 1998; Zhang & Mészáros, 2001).
@@
From an observational point of view, a small fraction of short GRBs that have plateau (or internal plateau) emission in X-rays point towards a stable NS (or a SMNS) as the central engine of short GRBs (Dai & Lu, 1998; Zhang & Mészáros, 2001; Rowlinson et al., 2013, 2010; Lü et al., 2015, 2017). However, due to their short lifetimes, it is difficult to observe definite clues for the HMNS as the central engine of short GRBs. Recently, Chirenti et al. (2023) discovered kilohertz QPOs in two short GRBs with a confidence level of about $5\sigma$ from archival Burst And Transient Source Experiment (BATSE) data, and suggest that these are evidence for a shortly surviving HMNS before the final collapse into a BH.
@@
Motivated by the above-described search in the BATSE data for a QPO that is the signature of an HMNS, this paper presents a systematic analysis of the short-duration GRBs observed with Fermi/Gamma-ray Burst Monitor (GBM) since its inception in 2008 and investigates whether such QPOs are also hidden in the light curves of those short GRBs. We systematically searched more than 3674 GRBs observed by Fermi/GBM and found 605 short-duration GRBs. We focused on analysing the 605 short-duration GRBs and found that three short-duration GRBs (e.g. GRBs 120323A, 181222B, and 190606A) with a high-frequency quasi-periodic signal can be identified with a confidence level above 5.2 $\sigma$. Our data reduction and QPO search method are shown in Section 2. The observations and results of our QPO search for those three GRBs are presented in Section 3. Conclusions are drawn in Section 4 with some additional discussions.
@@ 2.1 Fermi/GBM data reduction
The Fermi satellite was launched in 2008 June and has been in operation for 16 yr. There are two instruments onboard the Fermi satellite. One is the GBM (Meegan et al., 2009), which has 12 sodium iodide (NaI) and two bismuth germanate scintillation detectors covering the 8 keV–40 MeV energy band. The other is the Large Area Telescope (Atwood et al., 2009), which has an energy coverage from 20 MeV to 300 GeV.
@@
We find that the $T_{\rm 90}$ of 605 GRBs out of 3674 GRBs are less than 2 s at the official Fermi website. Here, we focus on studying the 605 short-duration GRBs and extract the light curves of those short-duration GRBs from 12 NaI detectors by adopting a 128 $\mu$s time bin.
@@ 2.2 QPO search method in short GRB samples
The data that we adopted consist of a flux time series of 605 short GRBs, and it includes 12 simultaneous light curves from the 12 NaI detectors in the GBM TTE (time-tagged event) data for each short GRB. We extracted the NaI detector light curve with a time bin of 128 $\mu$s in the energy range of 50–900 keV. Each NaI detector points to a slightly different position on the sky, the light curves of each GRB are ranked in order of increasing angular distance between the GRB position and detector pointing. They are designated as $LC_{i}$ where $i=1,2...,12$ from closest to farthest. 12 summed light curves are constructed by successively adding one more light curve to the sum in order of the ranking. These are designated as
@@
These summed light curves are not statistically independent. Next, we applied to $LC_{\rm sum}(n)$ and cut down the light curves to a duration of 2.048 s, which consists of 16,000 bins. Our selection of the start and end times of the light curves can be expressed as
@@
Here, the $a_{\rm j}$ is the result of the DFT, and $N$ is the number of bins in a light curve. Each power spectrum $I_{\rm j}$ is fitted with a broken power law plus constant using the maximum likelihood method (Vaughan, 2010) by assuming that the powers are exponentially ($\chi^{2}_{v}$ v with 2 degrees of freedom) distributed. The broken power-law function can be expressed as
@@
where $\alpha_{1}$ and $\alpha_{2}$ are the exponential slopes, respectively, $j_{\rm b}$ is the break point between the two slopes, and $C$ is the white noise component of the power spectrum. The likelihood function is equivalent to minimizing the function $D$ which is expressed as
@@
where $R_{\rm k-max}$ is the largest value in the power spectrum $R_{\rm k}(j)$. Each GRB contains 12 light curves, and we search each power spectrum for 10 values of $k$ and, therefore, obtain 120 $R_{\rm k-max}$ values for each GRB. In order to evaluate whether or not $R_{\rm k-max}$ is generated by quasi-periodic signals, one can calculate the probability that $R_{\rm k-max}$ is not produced by noise. A simulation is performed generating $10^{10}$ power spectra from $10^{10}$ simulated Poisson noise light curves. For each simulated power spectrum, 10 $R_{\rm k-max}$ ($k$=1,…,10) values are generated by the method described above. To estimate a confidence level associated with a given $R_{\rm k-max}$, the number of times $R_{\rm k-max}$ that this $R_{\rm k-max}$ is exceeded in the $R_{\rm k-max}$ simulated power spectra is counted. The probability $G_{\rm k}$ of a given $R_{\rm k-max}$ value is then calculated as
@@
where $G_{\rm k}$ is the probability that $R_{\rm k-max}$ is not generated by random noise. Therefore, for each GRB, we can obtain 120 different $G_{\rm k}$ values. If there is a quasi-periodic signal in a GRB, the maximum $G_{\rm k}$ value is the value that is most likely to be produced by the quasi-periodic signal. This procedure can be represented as
@@
Here, $S_{j}$ is the best fit with a broken power-law function, and it is not a free parameter in the fits. $A$ is the power density of the Lorentzian at its center, $v$ is the central frequency of the Lorentzian, and $\Delta v$ is the full width at half-maximum of the Lorentzian. We adopt the central frequency $v$ as the representative frequency of the quasi-periodic signal. By considering the frequency of the quasi-periodic signal which is distributed within a range, we use the 90 per cent integral area of the $L_{j}$ function within the Fourier frequency as the representative frequency range of the quasi-periodic signal. The lower limit and upper limit of the frequency correspond to the integral area 5 and 95 per cent, respectively.
@@ 4 Conclusions
Based on the data analysis of our QPO search for those three short GRBs, we find that the duration of a detected QPO is shorter than that of the burst. For example, the start and end times of the burst for GRBs 120323A, 181222B, and 190606A are [0, 0.384] s, [0.032, 0.576] s, and [0, 0.224] s, respectively. However, the start and end time of the detected QPO for those three short GRBs are [0.0384$\pm$0.0256, 0.0896$\pm$0.0256] s, [0.064$\pm$0.0256, 0.1664$\pm$0.0256] s, and [0.01$\pm$0.0256, 0.112$\pm$0.0256] s, respectively. On the other hand, the light curves of those three short GRBs seem to be composed of two pulses, and it is found that the QPO signals exist during the first pulse and the rising segment of the second pulse, but disappear during the decay segment of the second pulse. In general, two effects may result in such a situation. One is incorrect data analysis with the selected time-step, but we adopted a time-step of duration 0.0512 s which is much less than the burst duration. This smaller time-step is not likely to result in such a situation. The QPO signal disappearing during the decay segment of the second pulse may arise in the intrinsic physics, but this remains debatable.
@@
However, invoking an HMNS to interpret the high-frequency QPO in the prompt emission of those three short GRBs (e.g. GRBs 120323A, 181222B, and 190606A) imposes two curious questions that are not clarified in either theory or numerical relativity simulations of NS mergers. One question regards the composition and structure of an HMNS. Initially, the temperature should be high enough shortly after the formation of the HMNS, and the composition of the HMNS should be dominated by fluid instead of rigid structure because there has not been enough time to cool down. If this is the case, the oscillation of fluid seems to have difficulty producing such high-frequency quasi-periodic signals. The other question is how to produce the QPO signals only in the first pulse and the rising segment of the second pulse. There remains a poor understanding of this question.
In order to investigate the remnants of those three short GRBs-like events and find evidence for an HMNS in a double NS merger, more information from the simultaneous detection of both gravitational wave and multi-band electromagnetic radiation is essential. We therefore strongly encourage follow-up observations when the LIGO Virgo–KAGRA (LVK) Collaboration detects the gravitational wave signal of those three short GRBs-like events in the future.